Let n be an integer. Use Definition 1.6 to explain why 2n + 9

 is an odd integer 2n + 9 = 2(    )+ 1


1
Expert's answer
2021-03-31T15:03:09-0400

"2n+9" is odd because

"2n+9=2n+2*4+1=2(n+4)+1"